OpenStudy (anonymous):

use the substitution method to solve the following system of equations 2x y = 5 3x – 2y = 25

OpenStudy (andriod09):

is this it?

OpenStudy (anonymous):

yep

OpenStudy (andriod09):

what is the 2x y=5?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

srry

OpenStudy (andriod09):

thanks. so then it would be like this:

OpenStudy (andriod09):

\[2x+y = 5\] \[3x – 2y = 25\]

OpenStudy (anonymous):

yeppers

OpenStudy (andriod09):

i say we find x first. you agree?

OpenStudy (andriod09):

well, no. we could just add the equations to gether, and then simplify. that would be easyest.

OpenStudy (anonymous):

k

OpenStudy (andriod09):

okay., so then what we do: \[2x+y=5\]\[3x–2y=25\] would equal \[5x-y-30\]

OpenStudy (andriod09):

so what do you want to find first?

OpenStudy (anonymous):

Nevermind I don't need help anymore

OpenStudy (andriod09):

oki. :/

OpenStudy (anonymous):

lol sorry

OpenStudy (andriod09):

i need a medal to survive though. :S

OpenStudy (anonymous):

lol